Determine how many 1.5 V AA zinc chloride batteries will be needed to supply the voltage for a remote control. 9V and will consume 3.2 W of power.
1
Expert's answer
2015-09-02T01:52:28-0400
Determine how many 1.5 V AA zinc chloride batteries will be needed to supply the voltage for a remote control. 9V and will consume 3.2 W of power.

Given:
Vb=1.5 [V]
P=3.2 [W]
V=9 [V]


Solution:
According Battery reference book

https://books.google.com.ua/books?id=q58IX4BM7-0C&pg=SA5-PA18&lpg=SA5-PA18&dq=current+of+1.5+V+AA+zinc+chloride+batteries&source=bl&ots=lRWjgcEg8y&sig=YEef68od7JrWduVATq8RHcyT9NI&hl=ru&sa=X&ved=0CEgQ6AEwA2oVChMI4vXTvNLXxwIVxpRyCh0usA4l#v=onepage&q=current%20of%201.5%20V%20AA%20zinc%20chloride%20batteries&f=false

we need 6 battery pack only

V=6*Vb = 6*1.5=9 V 0<Ib<1.5
P=I*V -> I=P/V = 3.2/6 = 0.53 A
Answer: 6 batteries